OpenStudy (anonymous):

x-intercept 3, y-intercept 3/5.....the slope of the intercept form for the line is ??

OpenStudy (zzr0ck3r):

so you have the points (3,0) and (0,3/5) can you figure out the slope?

OpenStudy (anonymous):

@zzr0ck3r No

OpenStudy (zzr0ck3r):

slope is (y_1-y_0)/(x_1 - x_0) for you y_1 = 3/5 and y_0 = 0 x_1 = 0 and x_0 = 3

OpenStudy (zzr0ck3r):

\[\frac{y_1-y_0}{x_1-x_0}\] \[y_1 = \frac{3}{5},y_0 = 0, x_1 = 0, x_0 = 3\]

OpenStudy (zzr0ck3r):

I cant make it any easier:)

OpenStudy (zzr0ck3r):

hold on we are not done, did you find the slope?

OpenStudy (anonymous):

nope

OpenStudy (zzr0ck3r):

why?

OpenStudy (anonymous):

i dont know how realtalk

OpenStudy (anonymous):

i suck at algebra

OpenStudy (zzr0ck3r):

\[\frac{3/5-0}{0-3}=\frac{3/5}{-3}=\frac{3}{5*-3}=-\frac{1}{5}\]

OpenStudy (zzr0ck3r):

y=mx+b so m=-1/5 y=-1/5x+b b is the y intercept which they gave you b = 3/5 so y=-1/5x+3/5

OpenStudy (anonymous):

thanks 4 ya help i appreciate it @zzr0ck3r

OpenStudy (zzr0ck3r):

np